# Sustainable Transparency on Recommender Systems: Bayesian Ranking of Images for Explainability

### v2

@@ 4.2 Compared Models
We compare BRIE against several baseline models from the State of the Art to evaluate its performance and effectiveness in explanation ranking tasks:
@@
    ELVis diez2020towards : Predicts the preference of a (user, photograph) pair applying a multi-layer perceptron $g$ to the concatenation of the user’s latent embedding $\mathbf{U}_{u}$ and the photograph’s latent embedding $\mathbf{V}_{p}^{\prime}$ as input. The estimated preference can be represented as $\hat{\mathbf{R}}_{up}=\sigma(g([\mathbf{U}_{u};\mathbf{V}_{p}^{\prime}])$.
@@
    MF-ELVis paz2022sustainable : MF-ELVis predicts the preference of a (user, photograph) pair using a dot product between the user’s latent embedding $\mathbf{U}_{u}$ and the photograph’s latent embedding $\mathbf{V}_{p}^{\prime}$. The estimated preference can be represented as $\hat{\mathbf{R}}_{up}=\sigma\langle\mathbf{U}_{u},\mathbf{V}_{p}^{\prime}\rangle$.
@@ 4.3.1 Performance Evaluation Metrics
These performance evaluation metrics provide a comprehensive assessment of the models’ effectiveness in explanation ranking tasks. MRecall@$k$ and MNDCG@$k$ are computed with $k=10$ for test cases where there are 10 or more photos to rank and the associated user has at least 10 photos in the training set, to ensure that we have sufficient per-user data to measure the performance of the models effectively. For MAUC all test cases and users and considered, to capture the overall discriminative power of the models.
@@ 4.3.2 Efficiency and Compactness Metrics
All training times and carbon emissions are tracked using Codecarbon schmidt2021codecarbon .
@@ 4.4.1 Framework and Machine Configuration
To ensure reproducibility and a fair evaluation, we implemented BRIE and re-implemented ELVis and MF-ELVis in an efficient Python framework using PyTorch-Lightning [^1] https://www.pytorchlightning.ai/. We make this framework available in a public repository [^2] https://github.com/Kominaru/BRIE.
@@ 4.4.2 Model Hyperparameters
Originally, authors of ELVis [^3] https://github.com/pablo-pnunez/ELVis selected model hyperparameters through a grid search in the validation sets of each dataset, identifying a specific configuration per dataset; this approach was also shared by MF-ELVis [^4] https://github.com/Kominaru/tfg-komi.
In contrast, model hyperparameters for BRIE were selected through one random search in the validation set of Barcelona, training 25 models inside the following hyperparameter space: latent factors ($d$) ranging from 4 to 1024, learning rate ($lr$) ranging from $10^{-3}$ to $10^{-5}$, and dropout ranging from 0 to 0.8. A fixed batch size of $2^{14}$ was used, and an early stopping policy was implemented with a patience $p$=5, a delta $\delta$=$10^{-3}$, and a maximum of 100 epochs while tracking validation MAUC during training, as suggested by the BPR publication rendle2012bpr . nd all model parameters were initialized using a Xavier distribution glorot2010understanding and optimized using Adam kingma2014adam .
After a hyperparameter search on Barcelona’s validation set, we trained on all datasets with the selected configuration, without early stopping. We chose hyperparameters ($d$=64, $lr$=$10^{-3}$, dropout=0.75, 15 epochs) optimizing efficiency and compactness while consistently outperforming the State of the Art, prioritizing sustainability over marginal performance gains.
@@ 5 Results and Discussion
In this section, we compare BRIE with state-of-the-art and baseline methods using the experimental setup described in Section 4. We evaluate the models’ performance using standard metrics, assess their efficiency and sustainability in terms of time and energy consumption during training and inference, and investigate the impact of the number of latent factors on model performance.
@@ 5.1 Performance comparison
Table 2 presents the performance of all models on the six datasets using three evaluation metrics (MRecall@10, MNDCG@10, MAUC). BRIE consistently outperforms other models in all metrics and datasets, indicating its superior ability to provide explanations with user-created images. On average, BRIE achieves a relative improvement of approximately 7.5% over the second-best method in terms of MAUC, the most comprehensive metric. This suggests that BRIE’s utilization of Bayesian Pairwise Ranking as a learning goal effectively captures nuanced preferences and interactions between users and photographs, and is as such a learning goal better suited for explanation ranking tasks than surrogate classification goals.
@@ 5.2 Effiency and sustainability comparison
[…]
@@
[…]
@@ 6 Conclusions
In this work, we have explored the generation of personalized explanations in recommendation systems, specifically focusing on explaining recommendations using user-created photographs. Our approach, BRIE, has shown promising results in addressing the limitations of existing models: by leveraging a more adequate learning goal based on Bayesian Pairwise Ranking (BPR) and a more efficient handling of training data, BRIE achieves performance superior to state-of-the-art approaches and demonstrates remarkable efficiency, emitting significantly fewer carbon emissions during training and inference. Moreover, BRIE maintains its effectiveness even with reduced computational resources and a drastic decrease in the total model size. Our findings highlight the need for improved models that excel in terms of performance without sacrificing efficiency and sustainability; BRIE represents a clear step forward in this direction.
@@
Moving forward, there are important avenues for future work. Firstly, we can explore methods for improved negative sample selection during training to enhance the model’s ability to discriminate between positive and negative instances; selecting only “reliable negatives” for training has the potential to lead to improved performance and more reliable explanations. Secondly, addressing the “cold start” problem is crucial, enabling BRIE to provide personalized explanations to new users efficiently. Techniques such as transfer learning can be explored for this purpose, as currently the system needs to be partially re-trained to provide explanations to new users. Lastly, ensuring the scalability of BRIE is important as recommendation systems deal with increasingly larger datasets and user bases, and the model sizes of BRIE and other models scale linearly with the number of unique users. Investigating methods to optimize the model’s scalability and resource utilization will enhance its practical applicability.
7 Acknowledgement
@@
This research project has been supported by the National Plan for Scientific and Technical Research and Innovation of the Spanish Government (Grant PID2019-109238GB-C22), by the Spanish Ministry of Science and Innovation (Grant FPU21/05783), and by the Xunta de Galicia (Grant ED431C 2022/44) with the European Union ERDF funds. CITIC, as Research Center accredited by Galician University System, is funded by “Consellería de Cultura, Educación e Universidade from Xunta de Galicia”, supported in an 80% through ERDF Operational Programme Galicia 2014-2020, and the remaining 20% by “Secretaría Xeral de Universidades” (Grant ED431G 2019/01).

### v3

@@ 4.2 Compared Models
We compare BRIE against the existing models for the task, as well as other baseline approaches, to evaluate its performance and effectiveness in visual-based explanation through user-uploaded images in Recommender Systems:
@@
    ELVis [18]: Predicts the preference of a (user, photograph) pair applying a multi-layer perceptron $g$ to the concatenation of the user’s latent embedding $\mathbf{U}_{u}$ and the photograph’s latent embedding $\mathbf{V}_{p}^{\prime}$ as input. The estimated preference can be represented as $\hat{\mathbf{R}}_{up}=\sigma(g([\mathbf{U}_{u};\mathbf{V}_{p}^{\prime}])$.
@@
    MF-ELVis [19]: MF-ELVis predicts the preference of a (user, photograph) pair using a dot product between the user’s latent embedding $\mathbf{U}_{u}$ and the photograph’s latent embedding $\mathbf{V}_{p}^{\prime}$. The estimated preference can be represented as $\hat{\mathbf{R}}_{up}=\sigma\langle\mathbf{U}_{u},\mathbf{V}_{p}^{\prime}\rangle$.
@@ 4.3.1 Performance Evaluation Metrics
As done by previous works, the computation of MRecall@$k$ and MNDCG@$k$ is performed with $k=10$, and considering only the test cases that fulfil two conditions: First, the test case must contain at least 10 images to rank (1 positive example and at least 9 negative examples); this removes trivial test cases since, with $k=10$, any case with $\leq 10$ images would trivially return Recall@10=1 and NDCG@10$>$0. Second, the author user of the positive example in the test case must have least 10 positive examples (uploaded photographs) in the training set; removing these test cases that correspond to users in near-cold-start scenarios is common in user-interaction data to give a more reliable measure of ranking performance [28, 29]. On the other hand, when computing MAUC all test cases and users and considered, to have a measure the overall discriminative power of the models.
@@ 4.3.2 Efficiency and Compactness Metrics
All training times and carbon emissions are tracked using Codecarbon [30].
@@ 4.4.1 Framework and Machine Configuration
For the sake of reproducibility and a fair evaluation, we implemented BRIE and re-implemented ELVis and MF-ELVis in a common Python framework using PyTorch-Lightning [^1] https://www.pytorchlightning.ai/; all models share the same infrastructure for data loading, computation of metrics, and any other processes not unique to the particular design of a specific model. We make this framework available in a public repository [^2] https://github.com/Kominaru/BRIE.
@@ 4.4.2 Model Hyperparameters
Originally, the authors of ELVis [^3] https://github.com/pablo-pnunez/ELVis [18] selected model hyperparameters through a grid search in the validation sets of each dataset, identifying a specific configuration per dataset; this approach was also shared by MF-ELVis [^4] https://github.com/Kominaru/tfg-komi [19]. In this work, we train and evaluate both models using the single best hyperparameters per dataset reported in their original publications.
Model hyperparameters for BRIE were selected through one random search in the validation set of Barcelona; this is the smallest available dataset that does not exhibit irregular or volatile performance evaluation, as found by Díez et al. [18]. We trained 25 models with random hyperparameters inside the following search space: latent factors ($d$) ranging from 4 to 1024, learning rate ($lr$) ranging from $10^{-3}$ to $10^{-5}$, and dropout ranging from 0 to 0.8. A fixed batch size of $2^{14}$ was used, and an early stopping policy was implemented with a patience $p$=5, a delta $\delta$=$10^{-3}$, and a maximum of 100 epochs while tracking validation MAUC during training, as suggested by the BPR publication [27]. All model parameters were initialized using a Xavier distribution [31] and optimized using Adam [32].
After a hyperparameter search on Barcelona’s validation set, we trained on all datasets with the selected configuration, without early stopping. We selected the final hyperparameter combination based on both effectiveness and efficiency: considering only the configurations that consistently outperformed the State of the Art, we selected the one with the lowest inference emissions ($d$=64, $lr$=$10^{-3}$, dropout=0.75, 15 epochs), i.e. the one with the lowest $d$ (intuitively, the inference emissions of BRIE monotonically increase/decrease with larger/smaller $d$ values).
@@ 5 Results and Discussion
In this section, we compare BRIE with the existing state-of-the-art methods of the task and other baseline approaches using the experimental setup described in Section 4. We evaluate the models’ performance using standard metrics, assess their efficiency and sustainability in terms of time and energy consumption during training and inference, and investigate the impact of the number of latent factors on model performance.
@@ 5.1 Performance comparison
Table 3 presents the performance of all models on the six datasets using three evaluation metrics (MRecall@10, MNDCG@10, MAUC). BRIE consistently outperforms ELVis and MF-ELVis, the other two available models for the task, in all metrics and datasets, indicating its superior ability to provide explanations with user-uploaded images. On average, BRIE achieves a relative improvement of approximately 7.5% over the second-best method in terms of MAUC, the most comprehensive metric. This suggests that BRIE’s utilization of Bayesian Pairwise Ranking as a learning goal effectively captures nuanced preferences and interactions between users and photographs, and is as such a learning goal better suited for explanation ranking tasks than surrogate classification goals.
@@ 5.2 Effiency and sustainability comparison
In this subsection we compare the efficiency and sustainability of BRIE, ELVis, and MF-ELVis by comparing the evolution of AUC versus training time and CO2 emissions for all models and datasets, as well as inference times and CO2 emissions. Both training and inference are performed using GPU acceleration, and the final models used for reporting the results in Table 3 are considered.
Figure 6 compares AUC evolution versus training time and CO2 emissions for the different models. At the end of the training process, BRIE emits approximately 75% less CO2 than ELVis and 50% less than MF-ELVis, the current most sustainable model. Moreover, BRIE requires around three times less training time than ELVis in all cities, tying MF-ELVis as the model with the fastest training. The figure also reveals an interesting insight: BRIE consistently outperforms ELVis and MF-ELVis no matter the training time and emitted CO2, i.e. for any given training time or CO2 emissions goal, BRIE can achieve higher performance compared to the other models. Alternatively, if a specific performance goal is set, BRIE can attain it faster and with lower CO2 emissions.
@@
With respect to inference, Figure 7 shows BRIE and MF-ELVis present no significant variation and are consistently around two times faster than ELVis. Furthermore, BRIE emits 50% less CO2 during inference compared to both MF-ELVis and ELVis.
Overall, these findings can be attributed to several factors. Firstly, BRIE’s more adequate learning strategy based on Bayesian Pairwise Ranking (BPR) and handling of training data enables it to achieve higher performance early in the training process. Secondly, the complexity and resource requirements of the use of an MLP as a learned similarity function in ELVis incur penalties in terms of training and inference CO2 emissions and time compared to BRIE’s fixed similarity function based on dot product.
@@ 6 Conclusions
In this work, we have explored the generation of personalized explanations in recommendation systems, specifically focusing on explaining recommendations using user-uploaded photographs. Our approach, BRIE, has shown promising results in addressing the limitations of existing models by leveraging a more adequate learning goal based on Bayesian Pairwise Ranking (BPR) and a more efficient handling of training data. In a state-of-the-art evaluation scheme comprising six real-world datasets of explainability of restaurant recommendations, BRIE achieves performance consistently superior to approaches of the State of The Art and demonstrates remarkable efficiency, emitting significantly fewer carbon emissions during training and inference. Moreover, BRIE maintains its effectiveness even with reduced computational resources and a drastic decrease in the total model size.
We believe that, with the development of a model like BRIE, we provide a double contribution to the broad field of ethical Machine Learning. First, our approach increases the effectiveness of a novel approximation of visual-based explainability, increasing the potential for transparency of systems that deal with user-item interaction data that contain user-images, such as Recommender Systems, e-commerce platforms, or social networks. This effective transparency is an increasingly necessary trait, promoting user trust and fairness and constituting a cornerstone of AI regulations such as the EU’s AI Act [7, 8]. Second, our method exhibits a remarkably strong computational efficiency and remains disentangled from the overarching AI system it explains: not only it doesn’t require the generation of images or auxiliary information, as it uses exclusively existing, user-uploaded images to explain outputs, but it also heavily improves the required time, CO2 emissions and model sizes required to achieve this approach when compared to state-of-the-art models. This quest for more sustainable Artificial Intelligence or Green AI is also a crucial aspect of AI-specific regulations like the aforementioned AI Act, which obliges high-impact general purpose models such as Recommender Systems to measure and minimize their environmental impact [7, 8]. This way, in the context of Recommender Systems, BRIE constitutes a double step forward: it both enhances the transparency of Recommender Systems, and reduces the environmental impact needed to do so.
@@
Moving forward, there are important avenues for future work. Firstly, we can explore Positive Unlabelled Learning methods for improved negative sample selection during training to enhance the model’s ability to discriminate between positive and negative instances; selecting only “reliable negatives” for training has the potential to lead to improved performance and more reliable explanations. Secondly, addressing the “cold start” problem is crucial, enabling BRIE to provide personalized explanations to new users efficiently. Techniques such as transfer learning can be explored for this purpose, as currently the system needs to be partially re-trained to provide explanations to new users. Lastly, ensuring the scalability of BRIE is important as recommendation systems deal with increasingly larger datasets and user bases, and the model sizes of BRIE and other models scale linearly with the number of unique users. Investigating methods to optimize the model’s scalability and resource utilization will enhance its practical applicability.
7 Acknowledgements
@@
This research work has been funded by MICIU/AEI/10.13039/501100011033 and ESF+ (grant FPU21/05783), ERDF A way of making Europe (PID2019-109238GB-C22) and ERDF/EU (PID2021-128045OA-I00), and by the Xunta de Galicia (Grant ED431C 2022/44) with the European Union ERDF funds. CITIC, as Research Center accredited by Galician University System, is funded by “Consellería de Cultura, Educación e Universidade from Xunta de Galicia”, supported in an 80% through ERDF Operational Programme Galicia 2014-2020, and the remaining 20% by “Secretaría Xeral de Universidades” (Grant ED431G 2019/01).
Appendix A On the use of Dropout regularizations in BRIE
In Section 3, we discussed how, to prevent a empirically observed overfitting to training data by BRIE, we decided to apply a dropout regularization to both $U_{u}$ and $V^{\prime}_{p}$, the latent embeddings, to stabilize the learning process and further increase the performance of the model. This overfitting, which appears despite using a low number of latent factors $d$, is most likely caused due to the better suitability of the training policy used by BRIE for the image explanations ranking task. Figure 9 shows the evolution of Validation AUC and train loss $\mathcal{L}_{BPR}$ of a fitting process in the dataset used for hyperparameter searching, Barcelona, for two versions of BRIE. The left results correspond to a BRIE that uses the best hyperparameter combinations found in Section 4 ($d$=64, lr=$10^{-3}$, dropout $p=0.75$, 15 epochs), while the right results correspond to a BRIE that removes the dropout layer (i.e. $p=0$), maintaining the rest of the configuration. As it can be observed, while using no dropout layer allows a much tighter fit to the training data. corresponding to a lower train loss, the final performance of BRIE trained with a high dropout is considerably higher, without sacrificing the sustainability of the training process.
Figure 9: Comparison of overfitting in the training process (based on the train loss and validation AUC) of two setups of BRIE. One (left) uses the best found hyperparameter combination described in Section 4, while the other (right) removes the dropout regularization from that same configuration.